OpenStudy (anonymous):

Help (:

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

This is the problem

OpenStudy (anonymous):

Do you understand it?

OpenStudy (texaschic101):

-7 1/8 - (-9 1/2) ---- get rid of parenthesis -7 1/8 + 9 1/2 --- turn to improper fractions - 57/8 + 19/2 --- common denominator is 8 -57/8 + 76/8 = 19/8 = 2 thats not right...let me try something else -7 1/8 - (-9 1/2) -7 1/8 + 9 1/2 first do whole numbers.. -7 + 9 = 2 now do fractions -1/8 + 1/2 = -1/8 + 4/8 = 3/8 making the answer : 2 3/8.....thats better

OpenStudy (anonymous):

O: i never understood something so fast

OpenStudy (texaschic101):

sorry it took so long....

OpenStudy (anonymous):

Thank you so much @texaschic101 no your fine you took your time to help me so take as long as you need (:

OpenStudy (texaschic101):

glad to help :)